# Capacity Approximations for Insertion Channels with Small Insertion Probabilities

### v3

@@ Abstract
Channels with synchronization errors, exhibiting deletion and insertion errors, find practical applications in DNA storage, data reconstruction, and various other domains. Presence of insertions and deletions render the channel with memory, complicating capacity analysis. For instance, despite the formulation of an independent and identically distributed (i.i.d.) deletion channel more than fifty years ago, and proof that the channel is information stable, hence its Shannon capacity exists, calculation of the capacity remained elusive. However, a relatively recent result establishes the capacity of the deletion channel in the asymptotic regime of small deletion probabilities by computing the dominant terms of the capacity expansion. This paper extends that result to binary insertion channels, determining the dominant terms of the channel capacity for small insertion probabilities and establishing capacity in this asymptotic regime. Specifically, we consider two i.i.d. insertion channel models: insertion channel with possible random bit insertions after every transmitted bit and the Gallager insertion model, for which a bit is replaced by two random bits with a certain probability. To prove our results, we build on methods used for the deletion channel, employing Bernoulli$(1/2)$ inputs for achievability and coupling this with a converse using stationary and ergodic processes as inputs, and show that the channel capacity differs only in the higher order terms from the achievable rates with i.i.d. inputs. The results, for instance, show that the capacity of the random insertion channel is higher than that of the Gallager insertion channel, and quantifies the difference in the asymptotic regime.
Index Terms:
  Channels with synchronization errors, insertion channel, Gallager insertion channel, channel capacity, information stability, deletion channel.
@@ I Introduction
Synchronization-error channels, including deletion, repetition, and insertion channels, serve as effective models for a variety of realistic and practical scenarios. Such channel models find extensive applications in DNA coding [1, 2], DNA-based storage systems [3, 4], magnetic data storage systems [5], data reconstruction [6], and numerous other domains. Despite the wide range of application areas, the capacity of channels with synchronization errors remains an underexplored research area. In [7], Dobrushin establishes counterparts to Shannon’s discrete memoryless channel’s (DMC’s) capacity formulation [8] for a wide range of channels with synchronization errors. He demonstrates that the channel capacity for such channels can be defined by the maximal mutual information per bit and can be achieved through a random coding scheme. In [9, 10], the authors consider channels with synchronization errors governed by a Markov chain. It is shown that if the underlying Markov chain is stationary and ergodic, the information capacity exists and is equal to the coding capacity. This result generalizes Dobrushin’s coding theorem [7] on the information stability of channels with memoryless synchronization errors.
@@
Despite the early result establishing the information stability of the memoryless synchronization error channels, the calculation of the exact capacity proved extremely challenging. For instance, even for the seemingly simple channel model of independent and identically distributed (i.i.d.) deletion channel, the channel capacity is unknown. As such the literature on channels with synchronization errors predominantly focuses on establishing lower and upper bounds for channel capacity. In [11], Gallager proposed using a pseudo-random string to convolutional codes with sequential decoding to correct synchronization and substitution errors. Tree codes are utilized in [12] to obtain lower bounds for insertion/deletion channels. In [13], a lower bound for the deletion channel is established by demonstrating a connection between deletion channels and the Poisson-repeat channel using jig-saw decoding. This lower bound is further improved in [14] for the capacity of channels with independent and identically distributed (i.i.d.) deletions and duplications.
@@
To develop upper bounds on the capacity of channels with synchronization errors, typical technique is to employ a genie-aided approach with additional side information, which results in more tractable channel models. In [15], an upper bound on the deletion channel capacity is derived using undeletable markers. Similarly, Ref. [16] provides deletion channel upper bounds coupled with lower bounds, by using different genie-aided systems. In [17], an upper bound for the non-binary deletion channel is obtained. Furthermore, in [18], capacity upper bounds are obtained for more general synchronization error channels exhibiting insertions, deletions and substitutions. In [19], using a systematic approach based on convex programming and real analysis, the author derives upper bounds on the capacity of the binary deletion channel, channels with i.i.d. insertions and deletions, and the Poisson-repeat channel. This framework can be applied to obtain capacity upper bounds for any repetition distribution. A recent and comprehensive overview of the capacity results of synchronization-error channels is provided in [20].
@@
Parallel to the results establishing upper and lower bounds on the channel capacity, there have also been efforts determining the behavior of the capacity of the deletion channel in the asymptotic regime of small deletion probabilities. For instance, [21] demonstrates that the capacity of the deletion channel is comparable to that of the binary symmetric error channel in the asymptotic case for small deletion probabilities. The authors establish that the trivial capacity lower bound for this channel, given by $1-H(p)$, is tight in the limit. They provide an upper bound of $1-(1-o(1))H(p)$, where the $o(1)$ term vanishes as the deletion probability $p\rightarrow 0$. More in line with our work in this paper, the works presented in [22, 23] leverage Dobrushin’s coding theorem [7] to approximate the capacity of the binary deletion channel using series expansions up to the second and third leading terms, respectively, for small deletion probabilities. These studies establish the channel’s capacity in the asymptotic regime of small deletion probabilities by calculating the dominant terms of the capacity expansion. Additionally, they demonstrate that the optimal input distribution can be obtained by perturbing the i.i.d. Bernoulli$(1/2)$ input distribution. In [24], a similar approach is applied to duplication channels, where the channel capacity is computed for small values of the duplication probability using a series expansion. Note that, although the duplication channel is closely related to the insertion channel and shares similarities with our channel model, it duplicates existing bits/symbols rather than inserting random bits/symbols, hence no runs are lost or no new runs are created, significantly simplifying the capacity calculation. Indeed, the capacity for the duplication channel can be computed with arbitrary prevision for any duplication probability [25].
@@
In this paper, we adopt a similar systematic approach to insertion channels, which serve as practical models for certain real-world communication schemes including DNA storage. We consider two fundamental binary insertion channel models: 1) the simple insertion channel, in which a random bit is inserted after each bit with a certain probability, and 2) the Gallager insertion channel, in which each bit is replaced with two random bits with a certain probability. In both cases, we focus on small insertion probabilities. We explore the first two terms in the series expansion of these channel models’ capacities and establish their achievability using i.i.d. Bernoulli$(1/2)$ input distribution, and establish a matching converse for the dominant (lower order) terms.
@@
[…]
@@
$G_{1}=-$ $\log(e)+\frac{1}{2}\sum_{l=1}^{\infty}2^{-l-1}l\log l+\frac{1}{2}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(b+1)2^{-a}2^{-b}h\left(\frac{1}{b+1}\right).$
@@
[…]
@@
where $G_{1}\approx 0.4901$, and $G_{2}\approx-0.5865$, respectively. Approximations obtained by omitting the terms with higher order for these channel models are illustrated in Fig. 1. As may be anticipated, the Gallager insertion channel replaces a bit with two random bits; hence, for a specific input, the ambiguity of the channel output is higher compared to the simple insertion channel model with random bit insertion (without replacement). Therefore, intuitively, one would expect a lower rate for the Gallager insertion channel, which is verified by the results in Fig. 1.
@@
Figure 1: Capacity approximation for both insertion channels.
@@ II Preliminaries
We consider insertion channel $W_{n}$ which takes $n$-bit binary vector $X^{n}$ as an input and outputs $Y(X^{n})$. We note that the inserted bits are independent of both other insertions and the original input bits. Specifically, we focus on the following two specific channel models in this paper:
1) Simple Insertion Channel: There is a random bit insertion after each bit with probability $\alpha$. We define the insertion pattern for this channel model by vector $A^{n}\in\{0,1\}^{n}$ whose $i$-th element is defined by
$A_{i}=\begin{cases}1,&\text{if there is an insertion after the }i\text{-th bit of }X^{n}\\ 0,&\text{otherwise}\end{cases}$
@@
2) Gallager Insertion Channel: Each bit is replaced with two independent random bits with probability $\alpha$ as described in [26]. The insertion pattern for this model is defined by $A^{n}\in\{0,1\}^{n}$ whose $i$-th element is
$A_{i}=\begin{cases}1,&\text{if the }i\text{-th bit of }X^{n}\text{ is replaced }\\ &\text{with two random bits}\\ 0,&\text{otherwise}\end{cases}$
for $i\in\{1,2,\cdots,n\}$. For this channel model, we define two insertion bit vectors as $B^{n}=(B^{1,n},B^{2,n})$ where $B^{j,n}\in\{0,1\}^{n}$ with $j\in\{1,2\}$ and their $i$-th element is defined similarly to that of the insertion channel model, i.e.,
@@
In [7], Dobrushin showed that the maximum rate for reliable communication can be formulated by the maximal mutual information per bit for memoryless channels with synchronization errors. It is also shown that information stability holds and Shannon capacity exists. This approach can be used to express the capacity of the insertion channels above, as stated in the following theorem.
@@
(Taken from [7]) For the memoryless channels with finite expectation of output vector length, consider the definition
@@
Furthermore, as stated in [7, Th. 5], the input $(X_{1},X_{2},\cdots,X_{n})$ can be considered as consecutive coordinates of a stationary ergodic process as in the following lemma.
@@
(Taken from [23]) With $X_{i}\in\{0,1\}$ and stationary and ergodic process $\mathbb{X}=\{X_{i}\}_{i\in\mathbb{Z}}$, the limit $I(\mathbb{X})=\lim_{n\rightarrow\infty}\frac{1}{n}I(X^{n};Y(X^{n}))$ exists and we have
@@
See [23] for the proofs of Theorem 1 and Lemma 2.
@@
Throughout this paper, we will use the following definition: for the binary process $\mathbb{X}$, the consecutive blocks of 0’s and 1’s in $X=(X_{1},X_{2},\cdots,X_{n})$ are called runs, and we investigate the capacity of insertion channels with the stationary ergodic input process $\mathbb{X}$, denoted by $\mathcal{S}$.
@@ III Decomposing Rate: Understanding $I(X^{n};Y(X^{n}))$
In this section, we identify the different terms necessary to determine the maximum rate for both insertion channel models under consideration. First, we define the runs for the input bit vector, which consists of consecutive bits with the same polarity, and a corresponding vector representing the lengths of the output bits (including the inserted bits). Using the channel models, insertion pattern, and inserted bit vector defined in Section II, we analyze the contributing factors to the capacity of the simple and Gallager insertion channels, which will be computed in detail in subsequent sections.
We use $X(j)$ to represent the $j$-th run in the input vector $X^{n}$, while $Y(j)$ denotes the corresponding sequence of bits in the output vector $Y^{n^{\prime}}$. We define $K$ as the vector of $|Y(j)|$’s, where ||$\cdot$|| ⋅ | indicates the number of bits in $Y(j)$. Assuming that there are $M$ runs in $X(j)$, one can express this as:
@@
$H(Y,K|X^{n})$ =$\stackrel{{\scriptstyle r_{2}}}{{=}}H(Y,K,A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
=$\stackrel{{\scriptstyle r_{3}}}{{=}}H(Y,K|X^{n},A^{n},B^{n})+H(A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
=$\stackrel{{\scriptstyle r_{4}}}{{=}}H(A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
=$\stackrel{{\scriptstyle r_{5}}}{{=}}H(A^{n},B^{n})-H(A^{n},B^{n}|X^{n},Y,K),$ (13)

### v4

@@ Abstract
Channels with synchronization errors, exhibiting deletion and insertion errors, find practical applications in DNA storage, data reconstruction, and various other domains. The presence of insertions and deletions renders the channel with memory, complicating capacity analysis. For instance, despite the formulation of an independent and identically distributed (i.i.d.) deletion channel more than fifty years ago, and proof that the channel is information stable, hence its Shannon capacity exists, calculation of the capacity remained elusive. However, a relatively recent result establishes the capacity of the deletion channel in the asymptotic regime of small deletion probabilities by computing the dominant terms of its capacity expansion. This paper extends that result to binary insertion channels, determining the dominant terms of the channel capacity for small insertion probabilities and establishing capacity in this asymptotic regime. Specifically, we consider two i.i.d. insertion channel models: the simple insertion channel, where a random bit may be inserted after each transmitted bit, and the Gallager insertion model, for which a bit is replaced by two random bits with a certain probability. To prove our results, we build on methods used for the deletion channel, employing Bernoulli(1/2)(1/2) inputs for achievability and coupling this with a converse using stationary and ergodic input processes, and show that the channel capacity differs only in the higher order terms from the achievable rates with i.i.d. inputs. The results, for instance, show that the capacity of the simple insertion channel is higher than that of the Gallager insertion channel, and quantify the difference in the asymptotic regime.
@@ I Introduction
Synchronization-error channels, including deletion, repetition, and insertion channels, serve as effective models for a variety of realistic and practical scenarios. Such channel models find extensive applications in DNA coding [abroshan2019coding, balado2010embedding], DNA-based storage systems [heckel2019characterization, lenz2019coding], magnetic data storage systems [guan2014coding], data reconstruction [levenshtein2001efficient], and numerous other domains. Despite the wide range of application areas, the capacity of channels with synchronization errors remains an underexplored research area. In [dobrushin1967shannon], Dobrushin establishes counterparts to Shannon’s discrete memoryless channel’s (DMC’s) capacity formulation [shannon] for a wide range of channels with synchronization errors. He demonstrates that the channel capacity for such channels can be defined by the maximal mutual information per bit and can be achieved through a random coding scheme. In [10619598, morozov2024shannon], the authors consider channels with synchronization errors governed by a Markov chain. It is shown that if the underlying Markov chain is stationary and ergodic, the information capacity exists and is equal to the coding capacity. This result generalizes Dobrushin’s coding theorem [dobrushin1967shannon] on the information stability of channels with memoryless synchronization errors.
@@
Despite the early result establishing the information stability of the memoryless synchronization error channels, the calculation of the exact capacity proved extremely challenging. For instance, even for the seemingly simple channel model of independent and identically distributed (i.i.d.) deletion channel, the channel capacity is unknown. As such the literature on channels with synchronization errors predominantly focuses on establishing lower and upper bounds for channel capacity. In [gallager2000sequential], Gallager proposed using a pseudo-random string to convolutional codes with sequential decoding to correct synchronization and substitution errors. Tree codes are utilized in [zigangirov1969sequential] to obtain lower bounds for insertion/deletion channels. In [mitzenmacher2006simple], a lower bound for the deletion channel is established by demonstrating a connection between deletion channels and the Poisson-repeat channel using jig-saw decoding. This lower bound is further improved in [kirsch2009directly] for the capacity of channels with independent and identically distributed (i.i.d.) deletions and duplications.
@@
To develop upper bounds on the capacity of channels with synchronization errors, typical technique is to employ a genie-aided approach with additional side information, which results in more tractable channel models. In [diggavi2007capacity], an upper bound on the deletion channel capacity is derived using undeletable markers. Similarly, Ref. [fertonani2010novel] provides deletion channel upper bounds coupled with lower bounds, by using different genie-aided systems. In [rahmati2013upper], an upper bound for the non-binary deletion channel is derived. Furthermore, in [fertonani2010bounds], capacity upper bounds are obtained for more general synchronization error channels exhibiting insertions, deletions and substitutions. In [10.1145/3281275], using a systematic approach based on convex programming and real analysis, the author derives upper bounds on the capacity of the binary deletion channel, channels with i.i.d. insertions and deletions, and the Poisson-repeat channel. This framework can be applied to obtain capacity upper bounds for any repetition distribution. A recent and comprehensive overview of the capacity results of synchronization-error channels is provided in [cheraghchi2020overview].
@@
Parallel to the results establishing upper and lower bounds on the channel capacity, there have also been efforts determining the behavior of the capacity of the deletion channel in the asymptotic regime of small deletion probabilities. For instance, [5513746] demonstrates that the capacity of the deletion channel is comparable to that of the binary symmetric error channel in the asymptotic case for small deletion probabilities. The authors establish that the trivial capacity lower bound for this channel, given by 1H(p)1-H(p), is tight in the limit. They provide an upper bound of 1(1o(1))H(p)1-(1-o(1))H(p), where the $o(1)$ term vanishes as the deletion probability $p\rightarrow 0$. More in line with our work in this paper, the works presented in [kanoria2010deletion, kanoria2013optimal] leverage Dobrushin’s coding theorem [dobrushin1967shannon] to approximate the capacity of the binary deletion channel using series expansions up to the second and third leading terms, respectively, for small deletion probabilities. These studies establish the channel’s capacity in the asymptotic regime of small deletion probabilities by calculating the dominant terms of the capacity expansion. Additionally, they demonstrate that the optimal input distribution can be obtained by perturbing the i.i.d. Bernoulli(1/2)(1/2) input distribution. In [6457365], a similar approach is applied to duplication channels, where the channel capacity is computed for small values of the duplication probability using a series expansion. Note that although the duplication channel is closely related to the insertion channel and shares similarities with our channel model, it duplicates existing bits/symbols rather than inserting random bits/symbols, hence no runs are lost and no new runs are created, significantly simplifying the capacity calculation. Indeed, the capacity of the duplication channel can be computed with arbitrary precision for any duplication probability [4418490].
@@
In this paper, we adopt a systematic approach similar to that of [kanoria2010deletion, kanoria2013optimal], but applied to insertion channels, which serve as practical models for real-world communication scenarios such as DNA storage, where insertions are a prevalent type of error. We focus on two fundamental binary insertion channel models: 1) the simple insertion channel, in which each transmitted bit is followed by a random bit with a certain small probability, and 2) the Gallager insertion channel, in which each bit is replaced by two random bits with a certain small probability. In both cases, we focus on small insertion probabilities. We derive the first two terms in the asymptotic expansion of the channel capacities and establish their achievability using an i.i.d. Bernoulli(1/2)(1/2) input distribution. We also provide a matching converse for the dominant (lower-order) term, thereby obtaining an asymptotic approximation to the channel capacity, rather than a lower or upper bound.
While our analytical strategy draws inspiration from the techniques developed for the deletion channel [kanoria2010deletion, kanoria2013optimal], the insertion channel presents fundamentally different challenges. Notably, the statistical behavior of insertion errors complicates the alignment between input and output sequences in a manner distinct from deletions. As a result, the entropy calculations, derivation of bounds, and design of auxiliary distributions require novel adaptations, even when the high-level proof structure resembles prior work. As the first work that provides capacity approximations for the insertion channels, the results provide a significant step toward understanding the capacity of synchronization channels beyond deletions.
@@
$C_{1}(\alpha)=1+\alpha\log(\alpha)+G_{1}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon}),$ (1)
@@
$G_{1}=$- $\log(e)+\frac{1}{2}\sum_{l=1}^{\infty}2^{-l-1}l\log l+\frac{1}{2}\sum_{a=1}^{\infty}(a+1)2^{-a}h\left(\frac{1}{a+1}\right).$
@@
$C_{2}(\alpha)=1+\alpha\log(\alpha)+G_{2}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon}),$ (2)
@@
where $G_{1}\approx 0.4901$, and $G_{2}\approx-0.5865$, respectively.
@@
Figure 1: Capacity approximations for both insertion channels.
[…]
(a) Upper and lower bounds for the Gallager insertion channel model.
[…]
(b) Upper bound for the simple insertion channel model.
Figure 2: Bounds for the capacity of the insertion channels.
Approximations obtained by omitting the terms with higher order for these channel models are illustrated in Fig. 1. As anticipated, the Gallager insertion channel replaces each bit with two random bits; hence, for a given input, the ambiguity in the channel output is higher compared to the simple insertion channel model with random bit insertion (without replacement). Therefore, intuitively, one would expect a lower rate for the Gallager insertion channel, which is verified by the results in Fig. 1. We note that these approximations are asymptotically exact as $\alpha\rightarrow 0$, and we expect that they remain as good approximations for small $\alpha$; however, there are no guarantees for any specific non-zero $\alpha$ value. Furthermore, we provide upper and lower bounds for the Gallager insertion channel in Fig. 2(a). For the first upper bound (UB), we consider a genie-aided system in which the insertion positions are revealed to the receiver as side information. The resulting channel model is equivalent to a binary erasure channel with erasure probability $\alpha$, whose capacity is $1-\alpha$. This can be considered an upper bound for the Gallager insertion channel due to the use of side information. For the second UB, we adopt the approach proposed in [fertonani2010bounds] for the channels with insertions, deletions and substitutions and apply it to the Gallager insertion channel. This approach is also genie-aided, but only partial side information is provided to the receiver by dividing the input sequence into length-$L$ subsequences and revealing the length of the corresponding output subsequences, i.e., the receiver knows exactly how many bits are inserted in each subsequence. Clearly, this subchannel is memoryless and has finite input/output alphabets, and hence its capacity can be numerically evaluated using the Blahut-Arimoto algorithm (BAA) by providing the probability of each input/output pair. In this way, an upper bound for the channel capacity is obtained. In Fig. 2(a), the corresponding UB is shown for $L=8$. Since the revealed side information is more limited in this case, the resulting UB is, as expected, tighter compared to the first one.
For one of the lower bounds, Gallager [gallager_key] considered the use of convolutional codes over channels with synchronization errors and derived an expression that represents an achievable rate (or equivalently, a capacity lower bound) for channels with insertion, deletion, and substitution errors. By equating the probabilities of error types other than insertion error to zero, we also provide this result as a lower bound in Fig. 2(a). For the second lower bound, we consider the bound given in [6516943], which provides an analytical lower bound on the capacity of the binary Gallager insertion channel by considering independent uniformly distributed inputs and computing lower bounds on the mutual information between the input and output sequences. This bound outperforms the one in [gallager_key] for a wide range of insertion probabilities. We note that our provided approximation fits well between the given lower and upper bounds for the Gallager insertion channel model, as expected.
For the simple insertion channel, we compute the upper bound using the same genie-aided partial side-information approach as for the Gallager insertion channel, combined with the Blahut–Arimoto algorithm. The result for $L=8$ is shown in Fig. 2(b).
@@ II Preliminaries
We consider an insertion channel $W_{n}$ which takes $n$-bit binary vector $X^{n}$ as an input and outputs $Y(X^{n})$. We note that the inserted bits are independent of both other insertions and the original input bits. Specifically, we focus on the following two specific channel models in this paper:
1) Simple Insertion Channel: There is a random bit insertion after each bit with probability $\alpha$. We define the insertion pattern for this channel model by the vector $A^{n}\in\{0,1\}^{n}$, whose $i$-th element is defined by
[…]
@@
2) Gallager Insertion Channel: Each bit is replaced by two independent random bits with probability $\alpha$, as described in [gallager_key]. The insertion pattern for this model is defined by $A^{n}\in\{0,1\}^{n}$ whose $i$-th element is
[…]
for $i\in\{1,2,\cdots,n\}$. For this channel model, we define two insertion bit vectors $B^{n}=(B^{1,n},B^{2,n})$ where $B^{j,n}\in\{0,1\}^{n}$ with $j\in\{1,2\}$ and their $i$-th element is defined similarly to that of the insertion channel model, i.e.,
@@
In [dobrushin1967shannon], Dobrushin showed that the maximum rate for reliable communication can be formulated by the maximal mutual information per bit for memoryless channels with synchronization errors. It is also shown that information stability holds and Shannon capacity exists. This approach can be used to express the capacity of the insertion channels above, as stated in the following theorem.
@@
(Taken from [dobrushin1967shannon]) For the memoryless channels with finite expectation of output vector length, consider the definition
@@
Furthermore, as stated in [dobrushin1967shannon, Th. 5], the input $(X_{1},X_{2},\cdots,X_{n})$ can be considered as consecutive coordinates of a stationary ergodic process as in the following lemma.
@@
(Taken from [kanoria2013optimal]) With $X_{i}\in\{0,1\}$ and stationary and ergodic process $\mathbb{X}=\{X_{i}\}_{i\in\mathbb{Z}}$, the limit $I(\mathbb{X})=\lim_{n\rightarrow\infty}\frac{1}{n}I(X^{n};Y(X^{n}))$ exists and we have
@@
See [kanoria2013optimal] for the proofs of Theorem 1 and Lemma 2.
@@
Throughout this paper, we will use the following definition: for the binary process $\mathbb{X}$, the consecutive blocks of 0’s and 1’s in $X=(X_{1},X_{2},\cdots,X_{n})$ are called runs, and we investigate the capacity of insertion channels with the stationary ergodic input process $\mathbb{X}$, denoted by $\mathcal{S}$. Furthermore, we denote by $L_{0}$ the length of the run in $\mathbb{X}$ that contains position 0, and by $L_{1},L_{2},\ldots$ the lengths of the subsequent consecutive runs. For $L\in\mathbb{N}$, we define
$\mathcal{S}_{L}=\{\mathbb{X}\in\mathcal{S}:\Pr(L_{0}>L)=0\},$
that is, the class of stationary and ergodic binary processes whose run lengths are almost surely bounded by $L$.
@@ III Decomposing Rate: Understanding $I(X^{n};Y(X^{n}))$
In this section, we identify the different terms necessary to determine the maximum rate for both insertion channel models under consideration. First, we define the runs for the input bit vector, which consists of consecutive bits with the same polarity, and a corresponding vector representing the lengths of the output segments (including the inserted bits). Using the channel models, insertion pattern, and inserted bit vector defined in Section II, we analyze the contributing factors to the capacity of the simple and Gallager insertion channels, which will be computed in detail in subsequent sections.
We use $X(j)$ to represent the $j$-th run in the input vector $X^{n}$, while $Y(j)$ denotes the corresponding sequence of bits in the output vector $Y^{n^{\prime}}$ including any inserted bits; i.e., $Y(j)$ consists of the bits originating from $X(j)$ together with any inserted bits if an insertion occurs within $X(j)$. For instance, consider an input spanning from the $j$-th to the $(j+2)$-th run:
[0 0 1 1 1 1 0 0 0].[0\ 0\ 1\ 1\ 1\ 1\ 0\ 0\ 0].
Clearly, we have
$X(j)$ =[0 $=[0\$0],
$X(j+1)$ =[1 1 1 $=[1\$1\ 1\ 1],
$X(j+2)$ =[0 0 $=[0\$0\ 0].
Assume that the insertion channel introduces a ‘0’ after the first and third bits, and a ‘1’ after the eighth bit. With this insertion pattern, the resulting output runs, defined according to our convention, are as follows:
$Y(j)$ =[0 0 $=[0\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{0}}\$0],
$Y(j+1)$ =[1 0 1 1 $=[1\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{0}}\$1\ 1\ 1],
$Y(j+2)$ =[0 0 1 $=[0\$${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{1}}\ 0$].
Note that, as in this example, under our output run definition, the output runs may contain bits with different polarities.
We define $K$ as the vector of $|Y(j)|$’s, where $|\cdot|$ indicates the number of bits in $Y(j)$. Assuming that there are $M$ runs in $X(j)$, one can express this as:
@@
$H(Y,K|X^{n})$ $\stackrel{{\scriptstyle(r_{2})}}{{=}}H(Y,K,A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
$\stackrel{{\scriptstyle(r_{3})}}{{=}}H(Y,K|X^{n},A^{n},B^{n})+H(A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
$\stackrel{{\scriptstyle(r_{4})}}{{=}}H(A^{n},B^{n}|X^{n})-H(A^{n},B^{n}|X^{n},Y,K)$
$\stackrel{{\scriptstyle(r_{5})}}{{=}}H(A^{n},B^{n})-H(A^{n},B^{n}|X^{n},Y,K),$ (12)